Balance equation of : 20cm^3 of 0.5 mol dm^-3 hydrocloric acid + excess of magnesium powder and 20cm^3 of 0.5 mol dm^-3 sulphric acid + excess of magnesium powder.
1
Expert's answer
2020-11-26T05:09:26-0500

2HCl+Mg=MgCl2+H2

H2SO4+Mg=MgSO4+H2